# URL: ]https://photox.vscht.cz/URL: ]https://www.controlled-molecule-imaging.org

### v1

@@ Abstract
Aqueous radiation chemistry emerges through ultrafast proton transfer and ion-radical formation with unexplored energy-redistribution dynamics steering the subsequent reactions. We performed time-resolved disruptive probing on pure water dimer, $(\text{H}_{2}\text{O})_{2}$, to disentangle the post-ionization reactions. Through kinetic-energy-resolved ion imaging, we unraveled the dynamics in the $(\text{H}_{2}\text{O})_{2}^{+}$ ground state: at low-energy ($\mathord{\sim}$0.05 eV) ultrafast proton transfer ($\mathord{\sim}$19 fs) is followed by $\text{H}_{3}\text{O}^{+}+\text{OH}$ fragmentation ($\mathord{\sim}$360 fs). At higher energies, proton transfer becomes hindered ($\mathord{\sim}$60 fs) while the subsequent fragmentation becomes faster ($\mathord{\sim}$210 fs), evolving into coupled dynamics (>0.15>0.15 eV, $\mathord{\sim}$100 fs). Moreover, we observed $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization proceeding through a Zundel-like structure. This reveals how ion-radical formation in ionized hydrogen-bonded networks shapes reactivity in aqueous dynamics.
Radiation chemistry in aqueous environments is largely governed by water radiolysis [1, 2, 3]. Upon irradiation, water undergoes a cascade of reactions initiated by the formation of highly reactive ion–radical species [3, 4] that give rise to fundamental phenomena such as biological radiation damage, catalytic transformations, material corrosion, and radical-driven oxidation. A molecular-level characterization of transient reactive species formation and the associated energy redistribution is therefore essential for advancing technologies including radiation therapy [4, 5], wastewater treatment [6, 7], nuclear reactor coolant systems [4, 8], and spaceflight applications [9, 10].
@@
Figure 2: Recorded transient signals related to disruptive probing of the $\text{H}_{3}\text{O}^{+}$ channel and their analysis. A and B display transient signals of the $\text{H}_{3}\text{O}^{+}$ and $\text{H}_{2}\text{O}^{+}$ channels (blue dots and black error bars), respectively. Fits (red lines) were obtained using the parametric model described in Materials and Methods. The data (gray points) in the range (500,50)(-500,-50) fs were excluded from the fitting procedure due to probe-induced pre-excitation effects. The inset in A shows a velocity map of $\text{H}_{3}\text{O}^{+}$ with only the central part (yellow circle) originating from $(\text{H}_{2}\text{O})_{2}^{+}$ dynamics. C displays the result of simulations of $(\text{H}_{2}\text{O})_{2}^{+}$ dynamics in its ground electronic state ($\text{D}_{0}$), corresponding data for $\text{D}_{1}$$\text{D}_{3}$ are provided in Supplementary Text. D represents the background corrected delay- and energy-resolved $\text{H}_{3}\text{O}^{+}$ signal. E shows the comparison of the bleached (blue), pump-only (yellow), and simulated ($\text{D}_{0}$, red) kinetic energy distributions [21]. F displays the timescales of the $\text{H}_{3}\text{O}^{+}$ formation obtained from the kinetic energy-resolved analysis using the parametric models described in Materials and Methods; see also Supplementary Text for the individual fits.
Figure 3: Recorded transient signals related to disruptive probing of the $(\text{H}_{2}\text{O})_{2}^{+}$ channel. A and B display transient signals of $(\text{H}_{2}\text{O})_{2}^{+}$ and $\text{H}^{+}$, respectively. C shows the delay- and energy-resolved $\text{H}^{+}$ signals after subtraction of the pump-only contributions. D and E show transient $\text{H}^{+}$ signals for selected kinetic energy ranges. All fits (red lines) were obtained using the parametric model described in Materials and Methods. The data (gray points) in the range (500,50)(-500,-50) fs were excluded from the fitting procedure due to probe-induced pre-excitation effects.
@@
All acquired pump–probe signals were fitted using a parameterized model based on a linear combination of exponentially-modified-Gaussian (EMG) functions shown as red traces in Figure 2 and 3. Model details and extracted parameters for all measured transient signals are given in Materials and Methods and Supplementary Text. While the $(\text{H}_{2}\text{O})_{2}^{+}$, $\text{H}^{+}$, and $\text{O}^{+}$ traces are well described by a single-term EMG model, the other ion species, $\text{H}_{3}\text{O}^{+}$, $\text{H}_{2}\text{O}^{+}$, and $\text{OH}^{+}$, require a double-term EMG model. Three distinct timescales emerge in the acquired data: The first one represents an ultrafast process of 25…40 fs, corresponding to the drop in the $\text{H}_{3}\text{O}^{+}$ trace. A second timescale of several-hundred femtoseconds corresponds to the recovery of the $\text{H}_{3}\text{O}^{+}$ signal, i. e., 200…250 fs for $\text{H}_{3}\text{O}^{+}$, $\text{H}_{2}\text{O}^{+}$, and $\text{OH}^{+}$ and $\mathord{\sim}580$ fs for the $\text{O}^{+}$ transient. The third timescale, in the $(\text{H}_{2}\text{O})_{2}^{+}$ and $\text{H}^{+}$ signals, is on the order of 1 ps. The identification of the bleached $\text{H}_{3}\text{O}^{+}$ and $(\text{H}_{2}\text{O})_{2}^{+}$ channels as the sources of the observed transients ensures that the extracted timescales correspond directly to the $\text{H}_{3}\text{O}^{+}+\text{OH}$ fragmentation and $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization pathways. Thus, the PT and fragmentation dynamics leading to $\text{H}_{3}\text{O}^{+}$ formation are characterized by lifetimes of $39\pm 9$ fs and $250\pm 28$ fs. The $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization is largely completed at $\mathord{\sim}$1 ps.
@@
An inspection of Figure 2 D reveals distinct timescales associated with the individual kinetic-energy ranges. The timescales extracted for $0\ldots 0.3$ eV are summarized in Figure 2 F. This shows an increase of the fast-component timescale from 19 fs to 60 fs, associated with PT (orange), with a $\text{H}_{3}\text{O}^{+}$ kinetic-energy increase from 0 eV up to 0.15 eV. This slowing demonstrates the transition from a direct to a hindered PT process, indicating changes in the energy flow on the potential energy landscape. This is accompanied by a simultaneous decrease of the slow-component, fragmentation, timescale from 358 fs to 211 fs, associated with faster fragmentation at higher energies (blue). At higher kinetic energies, between 0.15 eV and 0.3 eV, the two timescale converge to $\mathord{\sim}100$ fs, obtained from a modified model assuming equal PT and fragmentation timescales. Thus, strikingly, the $\text{H}_{3}\text{O}^{+}$ formation dynamics exhibits a strong kinetic-energy dependence, with its characteristic timescale changing markedly across the released-energy range, see Figure 4 .
@@
We also observed a signal bleach for the $(\text{H}_{2}\text{O})_{2}^{+}$ transient signal shown in Figure 3 A. The parent ion $(\text{H}_{2}\text{O})_{2}^{+}$ does not have kinetic energy from fragmentation, preventing an kinetic-energy–resolved analysis. Nevertheless, the DP induced the transfer of $(\text{H}_{2}\text{O})_{2}^{+}$ population into the $\text{H}^{+}$ channel, Figure 3 B. These $\text{H}^{+}$ signals were analyzed in an energy-resolved manner analogous to the $\text{H}_{3}\text{O}^{+}$ case, providing the energy-resolved signals of $\text{H}^{+}$ shown in Figure 3 C. The time-dependent signals for the energy ranges $0\ldots 0.1$ eV and $1.0\ldots 1.5$ eV are shown in Figure 3 E. The $\text{H}^{+}$ signal in the $0\ldots 0.1$ eV range shows a rapid signal bleach in the first 100 fs, the origin of which remains unresolved. Most $\text{H}^{+}$ appear with kinetic energies of 0.3…1.5 eV, with a timescale of 0.70.71.41.4 ps. We ascribe the $\text{H}^{+}$ signal to the disruption of transient, non-stabilized $(\text{H}_{2}\text{O})_{2}^{+}$ and thus the observed picosecond timescale to $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization, as illustrated in Figure 4 .
@@ Acknowledgments
All data underlying the figures and all analysis scripts are provided as Supplementary Material. All processed data are publicly available at Zenodo (DOI: 10.5281/zenodo.18678891, link). All raw data are archived in DESY’s data repositories under ID 11019141 and are available upon reasonable request.
@@
Materials and Methods
Supplementary Text
Tutorial on disruptive probing

### v2

# arXiv:2602.17505v2[physics.chem-ph] 02 Mar 2026

@@ Abstract
Aqueous radiation chemistry emerges through ultrafast proton transfer and ion-radical formation with unexplored energy-redistribution dynamics steering the subsequent reactions. We performed time-resolved disruptive-probing experiment of pure water dimer, $(\text{H}_{2}\text{O})_{2}$, to disentangle the post-ionization reactions. Through kinetic-energy-resolved ion imaging, we unraveled the dynamics in the $(\text{H}_{2}\text{O})_{2}^{+}$ ground state: at low-energy ($\mathord{\sim}$0.05 eV) ultrafast proton transfer ($\mathord{\sim}$19 fs) is followed by $\text{H}_{3}\text{O}^{+}+\text{OH}$ fragmentation ($\mathord{\sim}$360 fs). At higher energies, proton transfer becomes hindered ($\mathord{\sim}$60 fs) while the subsequent fragmentation becomes faster ($\mathord{\sim}$210 fs), evolving into coupled dynamics (>0.15>0.15 eV, $\mathord{\sim}$100 fs). Moreover, we observed $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization proceeding through a Zundel-like structure. This reveals how ion-radical formation in ionized hydrogen-bonded networks shapes reactivity in aqueous dynamics.
Radiation chemistry in aqueous environments is largely governed by water radiolysis [1, 2]. Upon irradiation, water undergoes reaction cascades initiated by the formation of highly reactive ion–radical species [3, 4], which underpin biological radiation damage, catalysis, material degradation, and radical-driven oxidation [2]. A molecular-level characterization of the formation of transient reactive species and the associated energy redistribution is essential for advancing technologies including radiation therapy [2, 5], wastewater treatment [6, 7], nuclear–reactor-coolant systems [2, 8], and spaceflight applications [9, 10].
@@
Figure 2: Recorded transient signals related to disruptive probing of the $\text{H}_{3}\text{O}^{+}$ channel and their analysis. A and B display transient signals of the $\text{H}_{3}\text{O}^{+}$ and $\text{H}_{2}\text{O}^{+}$ channels (blue dots and black error bars), respectively. Fits (red lines) were obtained using the parametric model described in Materials and Methods SI1. The data (gray points) in the range (500,50)(-500,-50) fs were excluded from the fitting procedure due to probe-induced pre-excitation effects. The inset in A shows a velocity map of $\text{H}_{3}\text{O}^{+}$ with only the central part (yellow circle) originating from $(\text{H}_{2}\text{O})_{2}^{+}$ dynamics. C displays the result of simulations of $(\text{H}_{2}\text{O})_{2}^{+}$ dynamics in its ground electronic state ($\text{D}_{0}$), corresponding data for $\text{D}_{1}$$\text{D}_{3}$ are provided in Supplementary Text SI2. D represents the background corrected delay- and energy-resolved $\text{H}_{3}\text{O}^{+}$ signal. E shows the comparison of the bleached (blue), pump-only (yellow), and simulated ($\text{D}_{0}$, red) kinetic energy distributions [21]. F displays the timescales of the $\text{H}_{3}\text{O}^{+}$ formation obtained from the kinetic energy-resolved analysis using the parametric models described in Materials and Methods SI1; see also Supplementary Text SI2 for the individual fits.
Figure 3: Recorded transient signals related to disruptive probing of the $(\text{H}_{2}\text{O})_{2}^{+}$ channel. A and B display transient signals of $(\text{H}_{2}\text{O})_{2}^{+}$ and $\text{H}^{+}$, respectively. C shows the delay- and energy-resolved $\text{H}^{+}$ signals after subtraction of the pump-only contributions. D and E show transient $\text{H}^{+}$ signals for selected kinetic energy ranges. All fits (red lines) were obtained using the parametric model described in Materials and Methods SI1. The data (gray points) in the range (500,50)(-500,-50) fs were excluded from the fitting procedure due to probe-induced pre-excitation effects.
@@
All acquired pump–probe signals were fitted using a parameterized model based on a linear combination of exponentially-modified-Gaussian (EMG) functions shown as red traces in Figure 2 and 3. Model details and extracted parameters for all measured transient signals are given in Materials and Methods SI1 and Supplementary Text SI2. While the $(\text{H}_{2}\text{O})_{2}^{+}$, $\text{H}^{+}$, and $\text{O}^{+}$ traces are well described by a single-term EMG model, the other ion species, $\text{H}_{3}\text{O}^{+}$, $\text{H}_{2}\text{O}^{+}$, and $\text{OH}^{+}$, require a double-term EMG model. Three distinct timescales emerge in the acquired data: The first one represents an ultrafast process of 25…40 fs, corresponding to the drop in the $\text{H}_{3}\text{O}^{+}$ trace. A second timescale of several-hundred femtoseconds corresponds to the recovery of the $\text{H}_{3}\text{O}^{+}$ signal, i. e., 200…250 fs for $\text{H}_{3}\text{O}^{+}$, $\text{H}_{2}\text{O}^{+}$, and $\text{OH}^{+}$ and $\mathord{\sim}580$ fs for the $\text{O}^{+}$ transient. The third timescale, in the $(\text{H}_{2}\text{O})_{2}^{+}$ and $\text{H}^{+}$ signals, is on the order of 1 ps. The identification of the bleached $\text{H}_{3}\text{O}^{+}$ and $(\text{H}_{2}\text{O})_{2}^{+}$ channels as the sources of the observed transients ensures that the extracted timescales correspond directly to the $\text{H}_{3}\text{O}^{+}+\text{OH}$ fragmentation and $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization pathways. Thus, the PT and fragmentation dynamics leading to $\text{H}_{3}\text{O}^{+}$ formation are characterized by lifetimes of $39\pm 9$ fs and $250\pm 28$ fs. The $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization is largely completed at $\mathord{\sim}$1 ps.
@@
An inspection of Figure 2 D reveals distinct timescales associated with the individual kinetic-energy ranges. The timescales extracted for $0\ldots 0.3$ eV are summarized in Figure 2 F. This shows an increase of the fast-component timescale from 19 fs to 60 fs, associated with PT (orange), with a $\text{H}_{3}\text{O}^{+}$ kinetic-energy increase from 0 eV up to 0.15 eV. This slowing demonstrates the transition from a direct to a hindered PT process, indicating changes in the energy flow on the potential energy landscape. This is accompanied by a simultaneous decrease of the slow-component, fragmentation, timescale from 358 fs to 211 fs, associated with faster fragmentation at higher energies (blue). At higher kinetic energies, between 0.15 eV and 0.3 eV, the two timescales converge to $\mathord{\sim}100$ fs, obtained from a modified model assuming equal PT and fragmentation timescales. Thus, strikingly, the $\text{H}_{3}\text{O}^{+}$ formation dynamics exhibits a strong kinetic-energy dependence, with its characteristic timescale changing markedly across the released-energy range, see Figure 4 .
@@
We also observed a signal bleach for the $(\text{H}_{2}\text{O})_{2}^{+}$ transient signal shown in Figure 3 A. The parent ion $(\text{H}_{2}\text{O})_{2}^{+}$ does not have kinetic energy from fragmentation, preventing a kinetic-energy–resolved analysis. Nevertheless, the DP induced the transfer of $(\text{H}_{2}\text{O})_{2}^{+}$ population into the $\text{H}^{+}$ channel, Figure 3 B. These $\text{H}^{+}$ signals were analyzed in an energy-resolved manner analogous to the $\text{H}_{3}\text{O}^{+}$ case, providing the energy-resolved signals of $\text{H}^{+}$ shown in Figure 3 C. The time-dependent signals for the energy ranges $0\ldots 0.1$ eV and $1.0\ldots 1.5$ eV are shown in Figure 3 E. The $\text{H}^{+}$ signal in the $0\ldots 0.1$ eV range shows a rapid signal bleach in the first 100 fs, the origin of which remains unresolved. Most $\text{H}^{+}$ appear with kinetic energies of 0.3…1.5 eV, with a timescale of 0.70.71.41.4 ps. We ascribe the $\text{H}^{+}$ signal to the disruption of transient, non-stabilized $(\text{H}_{2}\text{O})_{2}^{+}$ and thus the observed picosecond timescale to $(\text{H}_{2}\text{O})_{2}^{+}$ stabilization, as illustrated in Figure 4 .
@@ Acknowledgments
All data underlying the figures and all analysis scripts are provided as Supplementary Material. All processed data are publicly available at Zenodo (DOI: 10.5281/zenodo.18678891, preview link). All raw data are archived in DESY’s data repositories under ID 11019141 and are available upon reasonable request.
@@
Materials and Methods SI1
Supplementary Text SI2
Tutorial on disruptive probing SI3